OpenStudy (anonymous):

Find the slope of the tangent line to the polar curve r=sin(4θ) at θ=π/8.

jimthompson5910 (jim_thompson5910):

this page may be of help http://tutorial.math.lamar.edu/Classes/CalcII/PolarTangents.aspx

OpenStudy (anonymous):

still am confused

jimthompson5910 (jim_thompson5910):

where are you stuck in the article?

OpenStudy (anonymous):

it has cos and ii am trying to find (sin) in my problem

jimthompson5910 (jim_thompson5910):

x = r*cos(theta) x = sin(4theta) * cos(theta) ... replace 'r' with 'sin(4theta)' what do you get when you derive both sides with respect to theta?

OpenStudy (anonymous):

r/r*cos ??

jimthompson5910 (jim_thompson5910):

if the theta's are throwing you off, then think of them as x ie think of it as y = sin(4x)*cos(x) what is dy/dx here?

OpenStudy (anonymous):

idk

jimthompson5910 (jim_thompson5910):

use the product and chain rule here